OpenStudy (anonymous):

Condense logm^50 + logm^12 a.logm600 b.logm(12 + 50) c.logm(50/12) d.log m50^12

OpenStudy (anonymous):

Log(m) + Log(n) = Log(mn) So, Log((m^50)*(m^12) = Log(m^62) is the answer...Looks like it is B..

jimthompson5910 (jim_thompson5910):

It's not B and stop giving out (the wrong) answers